L has y-intercept (0, -3) and is parallel to the line with the equation y=2/3x + 1,
L has y-intercept (0,2) and is perpendicular to the line with equation 2x - 3y=6.

Answer by stanbon(48568)

L has y-intercept (0, -3) and is parallel to the line with the equation y=2/3x + 1,
The slope is 2/3 and the y intercept is -3.
So the equation is y=(2/3)x-3
---------------
L has y-intercept (0,2) and is perpendicular to the line with equation
2x - 3y=6.
The slope of the given line is (2/3) so the slope of the perdicular line
must be -3/2
So the equation of the perpendicular line is y=(-3/2)x+2
==========
Cheers,
Stan H.